OpenStudy (priyar):

A coin is placed on a horizontal platform which undergoes vertical SHM of angular frequency "w". The amplitude of oscillation is gradually increased.the coin will leave contact with the platform for the first time at?

Parth (parthkohli):

\[\frac{N-mg}{m}=-\omega^2 x\]\[N\ge0 \Rightarrow \text{contact maintained}\]

OpenStudy (priyar):

at N=0 there is no contact..right?

Parth (parthkohli):

Yes.

OpenStudy (priyar):

then why did u include it in the abv. exp?

Parth (parthkohli):

Oh dunno why I typed \(\ge 0\).

Parth (parthkohli):

Yeah, it's just \(>0\).

OpenStudy (priyar):

ok...give me a min.. lemme try to do this

OpenStudy (priyar):

at x= g/w^2 right? btw where have used the fact that amplitude is gradually increasing?

Parth (parthkohli):

Nowhere, really. It's like saying we have an inclined slope with a gradually increasing angle, and it asks you to find the angle at which a block begins to slip.

ganeshie8 (ganeshie8):